OpenStudy (anonymous):

solve for the principal values of indicated variable: 2sin(2x)-sqrt3=0 and tan^(2) x-1=0

OpenStudy (anonymous):

ok: \[\sin(2x)=\frac{1}{2}\sqrt{3}\]

OpenStudy (anonymous):

When is the sine 1/2sqrt(3)?

OpenStudy (anonymous):

please clarify?

OpenStudy (anonymous):

For what value(s) of x does sin(x) equal \[\frac{1}{2}\sqrt{3}\]?

OpenStudy (anonymous):

60 degrees

OpenStudy (anonymous):

[pi/3]

OpenStudy (anonymous):

Is that the only solution?

OpenStudy (anonymous):

2pi/3

OpenStudy (anonymous):

right, additionally you can add multiples of 2pi to those solutions.

OpenStudy (anonymous):

so 2pi+(1/3)pi is also a solution.

OpenStudy (anonymous):

ok so what about the sin2x does the 2 not matter?

OpenStudy (anonymous):

it does matter, we're not done yet.

OpenStudy (anonymous):

2x=1/3pi, 2x= 2/3pi

OpenStudy (anonymous):

Let's be a bit more precise 2x=1/3pi +k*2pi OR 2x=2/3pi+k*2p for any constant k.

OpenStudy (anonymous):

got it thanks! can you help me with the second question:\[\tan ^{2}X -1=0\]?

OpenStudy (anonymous):

ok, \[\tan^2(x)=1\]

OpenStudy (anonymous):

ok what next sensei?

OpenStudy (anonymous):

we need to get rid of the square of course.

OpenStudy (anonymous):

do we do that by changing it to 1-cos2X ------- 1+cos2x or is there an easier way?

OpenStudy (anonymous):

tan^2(x)=tan(x)^2, so tan(x)=?

OpenStudy (anonymous):

wait, let's do this first: x^2=1, what's x?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

right, (-1)^2=1

OpenStudy (anonymous):

so what does tan(x) equal?

OpenStudy (anonymous):

1 or -1?

OpenStudy (anonymous):

right, so we have tan(x)=1 OR tan(x)=-1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when does tan(x) equal 1?

OpenStudy (anonymous):

pi/4 or 5pi/4

OpenStudy (anonymous):

or 9pi/4,or... this is the same as before, now you can add a multiple of pi to the answers.

OpenStudy (anonymous):

got it thanks!!

OpenStudy (anonymous):

You're welcome, you still need to do the tan(x)=-1 of course.

OpenStudy (anonymous):

yes 3pi/4 ect right?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

thank you